In a recent year, 6.5x10^8 metric tons of wheat were produced in the world. One metric ton is equivalent to 1000 kilograms. A grain of wheat weighs about .000008 kilograms. Find the number of grains of wheat that were produced in the world.

Sorry, i didn't know what category to place it in! But i think it has something to do with scientific notation. Thanks!! :D

Answer by jim_thompson5910(35256)

This is a unit conversion problem.

Given: 6.5 x 10^8 metric tons

Abbreviations for units:
KG = kilogram
MT = metric ton

Conversion factor: 1000 KG/1 MT


Multiply by the conversion factor to go from MT to KG


(6.5 x 10^8 MT) * (1000 KG/1 MT) = (6.5 x 10^8)*(1 x 10^3) = 6.5 x 10^11 KG

Notice how the MT units cancel

So, 6.5 x 10^8 MT = 6.5 x 10^11 KG

-------------------------------------------------------
Now use the fact that "A grain of wheat weighs about .000008 kilograms." to do another conversion

Conversion factor: 1 grain/0.000008 KG


Multiply by the conversion factor to go from KG to number of grains


(6.5 x 10^11 KG)*(1 grain/0.000008 KG)
(6.5 x 10^11 KG)*(1 grain/8*10^(-6) KG)
(6.5/8) x (10^11/10^(-6)) grains
(0.8125) x 10^(11-(-6)) grains
(0.8125) x 10^(11+6) grains
(8.125 x 10^(-1)) x 10^(17) grains
8.125 x 10^(-1+17) grains
8.125 x 10^(16) grains

The answer in scientific notation is 8.125 x 10^(16) grains of wheat

Expand that all out, to write into standard form, to get 81,250,000,000,000,000 grains (8125 followed by 13 zeros) of wheat

Note: In words, that large number above is eighty-one quadrillion, two hundred fifty trillion